OpenStudy (anonymous):

Find the largest interval where the function is changing as requested. Decreasing f(x)=-sqrt(x+3) @MIT 18.01 Single …

OpenStudy (anonymous):

The function is decreasing for (-∞,0)

jhonyy9 (jhonyy9):

the zero point will be x=-3

jhonyy9 (jhonyy9):

,,Ullere" why ??? when for x=6 ,f(x)= -3 so for x=13,f(x)=-4 so from this result that the function is decreasing too

OpenStudy (anonymous):

Ah, it is because I didn't read the question correctly, I missed the negative sign outside of the brackets. the function would be decreasing for (0,[\infty\])

OpenStudy (anonymous):

Wouldn't it be decreasing from [-3,-∞] then ?

OpenStudy (anonymous):

Ooops I meant [-3,∞]****

OpenStudy (anonymous):

\[f(x)=-(x+3)^{1/2}\] ; \[x \ge-3\] \[f'(x)=-(1/2)(x+3)^{-1/2}\] f'(x)=0 ; x=-3 Test Value: x=0 \[f'(0)<0\] so function is decreasing on interval \[(-3,\infty)\]

OpenStudy (anonymous):

is first derivative test

OpenStudy (anonymous):

I knew it! Thanks

OpenStudy (anonymous):

No problem

OpenStudy (anonymous):

\[(-3,\infty)\]